OpenStudy (skarlettrose186):

Will Medal!! Which of the following values would complete the ordered pair if the point is on the graph of ƒ(x) = -2x + 3? (-1, ___) -4 0 1 5

OpenStudy (ryandean123):

-4

OpenStudy (518nad):

when you plug in -1 for x, what is f(x) f(-1) = -2(-1)+3

OpenStudy (518nad):

f(-1) = -2(-1)+3 =2+3 =5

OpenStudy (ryandean123):

HMM i had that smae out comee